# Parallel Batch-Dynamic Maximal Matching with Constant Work per Update

### v1

@@ 4. Algorithm
We now describe our parallel batch-dynamic algorithm. Roughly, the algorithm follows the structure of prior level-based dynamic maximal matching algorithms (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024). In these algorithms, and ours, inserting edges and deleting unmatched edges is reasonably straightforward, and the difficulty comes in deleting matched edges. When deleting a matched edge we need to find if any of the incident edges can be matched. If the degree of the edge is high this can be costly. The idea of the prior algorithms is to randomly select a possible edge to match across a sufficiently large sample (roughly its degree). The oblivious adversary therefore does not know which of the sampled edges is selected and in expectation will delete half the edges before reaching the one that was selected. When a matched edge is encountered, the algorithms, roughly speaking, need to explore neighboring edges, but this cost is amortized against the prior cheap deletes of unmatched edges.
@@
Our algorithm most resembles Assadi and Solomon’s hyperedge algorithm (Assadi and Solomon, 2021), although, beyond being parallel, it differs in important ways that are described in Section 1.1.
@@
    Every matched edge $e$ that owns $s$ sample edges is assigned a level $l=\lfloor\lg s\rfloor$.
@@
To efficiently access a leveled matching structure we will use the following data structures.
@@
Table 1 summarizes the data structures and other notation. Figure 2 lists the pseudocode for our algorithm. The user-level functions are insertEdges and deleteEdges, which insert and delete, respectively, batches of edges. Four of of the functions are just for adding and removing matched and cross edges. They just update the relevant data structures to maintain the invariants. When adding a match we also add all its sample edges. When removing a match, we remove its cross edges and mark its vertices as free. Removing a match assumes the sample edges have already been converted to cross edges.
Inserting a batch of edges first identifies which of the edges are free (are not incident on a matched edge) and runs a maximal matching on these. This could be any maximal matching. It then adds the matched edges to the leveled matching structure, each with just themselves as the sample. This places them on level 00 since $\lg 1=0$. Any remaining edges are added as cross edges.
@@
The function deleteMatchedEdges first takes all the sample edges owned by the matched edges to be removed and inserts them as cross edges. We note that when inserting them as cross edges they might or might not be owned by the same matched edge. The algorithm then separates the matched edges into two kinds: light and heavy. The light matches are ones for which the number of cross edges it owns is not much larger than its sample size, and in particular for a match $e$, less than $4r^{2}2^{l(e)}$. We allow ourselves to charge work proportional to the sample size to each match, and hence can afford to rematch these owned edges directly. The algorithm therefore simply removes the light matched edges using removeMatch, which also removes all their owned cross edges. The algorithm then reinserts the cross edges with insertEdges. [^6] In Solomon’s algorithm this is referred to as deterministic settling. The insertEdges will find new matches for the removed cross edges, if possible.
We are now left with deleting the heavy matches. We cannot afford to find matches for the cross edges owned by the heavy matched edges without taking new samples, which is performed by random settling. Before random settling, we remove from the structure all the heavy matches along with their owned cross edges (using removeMatch). Random settling is then performed in rounds. One round, i.e., one call to randomSettle, first takes all cross edges owned by the heavy matches and runs a random greedy matching on them. The output of this matching will be a maximal set of matched edges, along with a sample set for each. Each matched edges $e$ is added to the data structure with their sample $S_{e}$ at the appropriate level ($\lfloor\lg|S_{e}|\rfloor$) using addMatch.
@@
At the end of a round of random settling we therefore have two sets of matched edges to be deleted: the bloated and the stolen edges. We delete these edges using deleteMatchedEdges. As already described, this will partition them into light and heavy [^7] The bloated edges will always be heavy. The light ones can be directly removed and their owned edges reinserted with insertEdges. The heavy ones are removed and their owned edges are returned to be matched in the next round of random settle.
@@ 5. Analysis
We now analyze the cost bounds. We consider both work and 0pt, although the 0pt analysis is straightforward and left until the end. As in prior dynamic maximal matching algorithms (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024), it is useful to consider the lifetime of a match from when it is created until it is deleted. This is referred to as an epoch. It is also useful to distinguish epochs that are deleted by the user in the call to deleteEdge, vs. ones that are deleted internally by the algorithm by a random settle. The first kind are referred to as natural deletions of an epoch and the second as induced deletions of an epoch. In the following we will equate the epoch $e$ with the edge $e$ corresponding to its match.
@@ 5.1. Direct Work: Charging to all Epochs
The expected work of insertEdges$(E)$ is $O(r|E|)$.
@@
The cost of each round of random settle is either covered by matches in the current round or matches in the next round. Specifically, all but raising of matches edges (in raiseMatches) that end up being bloated, and the cost of deleteMatchedEdges at the end, will be covered by the current round.
@@
Putting the work costs together we get the following bound on work.
@@
Follows from Lemma 5.2, 5.3, and 5.4, which cover all work. ∎
@@ 5.2. Total Work: Charging to Natural Epochs
The input edge set $E$ to randomSettle is the union of the owned cross edges of a set of heavy deleted matches $M_{\mbox{\footnotesize old}}$. Each newly matched edge $m\in M_{\mbox{\footnotesize new}}$ can therefore, for accounting purposes, be associated with the match in $M_{\mbox{\footnotesize old}}$ that previously owned it, and we will use $p(m)$ to indicate this deleted match. Each matched edge $m\in M_{\mbox{\footnotesize new}}$ can contribute up to $r-1$ stolen deletes, one for each of its vertices except at least one incident on $p(m)$. Now each $e\in M_{\mbox{\footnotesize old}}$ can have at most $r$ newly matched edges incident on it (otherwise it would not be a matching). Therefore we can associate $r(r-1)$ stolen deletes with each $e\in M_{\mbox{\footnotesize old}}$. Furthermore we also associate the one bloated delete that might have caused each $e\in M_{\mbox{\footnotesize old}}$ to rise, and hence there are fewer than $r^{2}$ deletes associated with deleted edge $e$. This covers all the deletes.
@@
$S_{a}\;\;=\;\;\sum_{e\in M_{\mbox{\footnotesize new}}}|S(e)|\;\;=\;\;\sum_{e\in M_{\mbox{\footnotesize old}}}|C(e)|\;\;\geq\;\;4r^{2}\sum_{e\in M_{\mbox{\footnotesize old}}}2^{l(e)}\;\;>\;\;2S_{d}$
We note that having the gap between levels be a constant factor ($\alpha=2$ in our case) instead of proportional to $r$ (as in Assadi and Solomon (Assadi and Solomon, 2021)) is important to the proof. In particular, in the last step of the proof of Lemma 5.6 we lose a factor of $\alpha$. Since we have $\alpha=2$, this brings the coefficient from $4$ to $2$ for us, which is fine. But if $\alpha$ was $\Theta(r)$, then the coefficient would be reduced from $4$ to $\Theta(4/r)$, which is too small to be useful. We could have compensated for $\alpha=\Theta(r)$ by increasing the cutoff for being heavy (in isHeavy) by a factor of $r$, to $\Theta(r^{3}2^{l(e)})$, but then the amortized cost of the overall algorithm would also have increased by a factor of $r$, to be $O(r^{4})$ per edge update. Thus, we save a factor of $r$ by using thinner levels. In AS, reducing the level gap from $\Theta(r)$ to $2$ is not necessary because they assign every new match to some specific deletes. This extra cost consideration in our algorithm comes from “shuffling” the sampled edges during the matching.
@@
Let $S_{a}$ be the total sample size of all new epochs, $S_{i}$ the total sample size of all induced deletions and $S_{n}$ of natural deletes. From Lemma 5.6 we have that in each round of random settle the number of added samples is at least twice the number of samples from induced deletions. At the end of a batch update, the algorithm runs a insertEdges on the remaining unmatched edges. These are therefore not covered by the sample set of a random settle. Due to the termination condition, however, we know that the sample size of previously matched edges is twice as large as the number of the remaining edges. Let $|bloated|$ be the sample size of bloated remaining edges. Because these edges are bloated, they accrued a factor of at least $2r^{2}$ cross edges after rising. Thus, the sample size of deleted matches from the last round of random settle is at most $\frac{|\mbox{bloated}|}{2r^{2}}\leq\frac{|\mbox{bloated}|}{2}$. Therefore, the total sample space of induced deletions is at most the $\frac{1}{4}S_{a}$ from the final round plus the $\frac{1}{2}S_{a}$ from all other rounds, giving $S_{a}>\frac{4}{3}S_{i}$ within each batch deletion and hence also across all batch deletions. Since we start and end empty, we have that $S_{a}=S_{i}+S_{n}$. Putting these together we get $S_{n}>\frac{1}{3}S_{i}$.
@@ 5.3. Overall Cost Bounds
We now consider the overall work and 0pt bounds. Firstly we note that if we have a bound for the work from a start where the graph is empty to an end when it is empty again, this implies a bound that is at least as good for any prefix at least half as long as the whole sequence. All runs are a prefix of at least half of an empty-to-empty sequence since we can just delete the edges and the number of final deletes is at most the number of prior operations. Hence we need only consider empty-to-empty runs to asymptotically bound the cost for the prefix.
Let $Q$ be the set of natural epochs. Let $d_{1},d_{2},\ldots,d_{T}$ be the set of deletions of sampled edges by the user. Note that different $d_{i}$’s may come from different instances of parallelGreedyMatch and that $T\leq N$. Suppose that we dynamically assign prices to sampled edges as described in Section 3.1.
@@
The expected total natural sample space created is no more than the expected total price of all user deletes: $\mathbb{E}[\sum_{e\in Q}|S_{e}|]\leq\mathbb{E}[\sum_{t=1}^{T}\Phi^{\prime}(d_{t})]$.
@@
Because we start and end with an empty graph, note that the sum of the sample space created for all natural epochs is the same as the sum of the sample space of all natural epochs that is destroyed. Note that a natural epoch $e$ is destroyed by a user deletion and contributes $|S_{e}|$ to the total natural sample space destroyed.
Consider the user delete of an edge $e$. If $e$ is a cross edge, deleting it destroys no sample space, so we do not need to associate it to an epoch. If the edge is sampled, associate its delete with the epoch $e^{\prime}$ that owns it. Note that a delete is sampled iff it is early. Note that $e$ became sampled during some instance of parallelGreedyMatch. By Lemma 3.1, a sampled edge is owned by exactly one epoch within the instance of parallelGreedyMatch that sampled it. Because sampled edges do not participate in the algorithm until their matched edge is deleted, this edge will not join the sample space of any other epochs. Thus, the sample spaces of different instances of parallelGreedyMatch do not interfere with each other.
Let $e^{\prime}$ be a natural epoch. Because we end on an empty graph, $e^{\prime}$ is eventually deleted. Then applying Lemma 3.5, we get the total price of sampled deletes on $e^{\prime}$ is $|S_{e^{\prime}}|$. Since for each natural epoch, the sample space lost is no more than the price paid, the lemma follows. Note that the total price paid is not strictly equal to the total natural sample space lost because unmatched sampled deletes from induced epochs pay price but do not destroy natural sample space. ∎
@@
By Lemma 5.8, observe that $\mathbb{E}[\sum_{e\in Q}|S_{e}|]\leq\mathbb{E}[\sum_{t=1}^{T}\Phi^{\prime}(d_{t})]$. By Lemma 3.4, observe that $\mathbb{E}[\sum_{t=1}^{T}\Phi^{\prime}(d_{t})]\leq\sum_{t=1}^{T}\mathbb{E}[\Phi(d_{t})]\leq\sum_{t=1}^{T}2=2T$. By Lemma 5.7, the total work of the algorithm is bounded by $\sum_{e\in Q}Cr^{3}2^{l(e)}\leq\sum_{e\in Q}2Cr^{3}|S_{e}|\leq 2Cr^{3}(2T)\leq O(r^{3}N)$, where $C$ is the constant hidden by Lemma 5.7. ∎

### v2

@@ 4. Algorithm
We now describe our parallel batch-dynamic algorithm. Roughly, our algorithm follows the structure of prior level-based dynamic maximal matching algorithms (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024). In these algorithms, and ours, inserting edges and deleting unmatched edges is reasonably straightforward, and the difficulty comes in deleting matched edges. When deleting a matched edge we need to find if any of the incident edges can be matched. If the degree of the edge is high this can be costly. The idea of the prior algorithms is to randomly select a possible edge to match across a sufficiently large sample (roughly its degree). The oblivious adversary therefore does not know which of the sampled edges is selected and in expectation will delete half the edges before reaching the one that was selected. When a matched edge is encountered, the algorithms, roughly speaking, need to explore neighboring edges, but this cost is amortized against the prior cheap deletes of unmatched edges.
@@
    When an edge $e$ becomes matched, let $s$ be the number of sample edges $e$ owns. We assign $e$ the level $l=\lfloor\lg s\rfloor$.
@@
To efficiently access the leveled matching structure we will use the following data structures.
@@
Table 1 summarizes the data structures and other notation. Figure 3 lists the pseudocode for our algorithm. The user-level functions are insertEdges and deleteEdges, which insert and delete, respectively, batches of edges. Four of the functions are just for adding and removing matched and cross edges. They just update the relevant data structures to maintain the invariants. When adding a match we also add all its sample edges. When removing a match, we release its cross edges and mark its vertices as free. Removing a match assumes the sample edges have already been converted to cross edges.
Note that we do not initialize $P(v,l)$ to empty for every level at the beginning of the algorithm (which would lead to $\Theta(n\log n)$ work as there are $O(\log n)$ levels). Instead, we store the id of each initialized bag in a hash table. Before accessing a bag, we check whether it has been initialized, and initialize it if it has not yet been initialized.
Before describing how we handle batch insertions and deletions, we remark that our algorithm is quite different from Ghaffari and Trygub (Ghaffari and Trygub, 2024). Some of the differences include:
  1. (1)
    Our leveling structure has a gap of 2, rather than $\Theta(r)$.
  2. (2)
    Our leveling structure permits matched edges that gain many cross edges to remain at a low level, instead of forcibly resettling at a higher level.
  3. (3)
    In our algorithm, edges take on all ownership responsibilities: (matched) edges, not vertices, own edges.
  4. (4)
    We process matched deletions at all levels at once, instead of correcting levels from high to low.
  5. (5)
    We use random static maximal matching to randomly select matched edges, instead of iteratively choosing independent sets of edges to match via coin flips with increasing probability of success.
Our algorithm most resembles Assadi and Solomon’s hyperedge algorithm (Assadi and Solomon, 2021), although, beyond being parallel, it differs in important ways. For an extended discussion of the differences between our algorithm and prior work, see Section 1.1.
Inserting a batch of edges first identifies which of the edges are free (are not incident on a matched edge) and runs a maximal matching on these. This could be any maximal matching. It then adds the matched edges to the leveled matching structure, each with just themselves as the sample. This places them on level 0 since $\lg 1=0$. Any remaining edges are added as cross edges.
@@
The function deleteMatchedEdges first takes all the sample edges owned by the matched edges to be removed and inserts them as cross edges. We note that when inserting them as cross edges they might or might not be owned by the same matched edge. The algorithm then separates the matched edges into two kinds: light and heavy. [^8] Note that we handle light and heavy edges differently for efficiency (our work bounds), not for correctness. For example, if we designated all edges to be light, the matching would still be maximal. The light matches are ones for which the number of cross edges it owns is not much larger than its sample size, and in particular for a match $e$, less than $4r^{2}2^{l(e)}$. We allow ourselves to charge work proportional to the sample size to each match, and hence can afford to rematch these owned edges directly. The algorithm therefore simply removes the light matched edges using removeMatch, which also removes all their owned cross edges. The algorithm then reinserts the cross edges with insertEdges. [^9] In Solomon’s algorithm this is referred to as deterministic settling. The insertEdges function will find new matches for the removed cross edges, if possible.
We are now left with deleting the heavy matches. We cannot afford to find matches for the cross edges owned by the heavy matched edges without taking new samples, which is performed by random settling. Before random settling, we remove from the structure all the heavy matches along with their owned cross edges (using removeMatch). Random settling is then performed in rounds. One round, i.e., one call to randomSettle, first takes all cross edges owned by the heavy matches and runs a random greedy matching (ParallelGreedyMatch) on them. The output of this matching will be a maximal set of matched edges, along with a sample set for each. Each matched edges $e$ is added to the data structure with their sample $S_{e}$ at the appropriate level ($\lfloor\lg|S_{e}|\rfloor$) using addMatch.
@@
At the end of a round of random settling we therefore have two sets of matched edges to be deleted: the bloated and the stolen edges. We delete these edges using deleteMatchedEdges. As already described, this will partition them into light and heavy. [^10] The bloated edges will always be heavy The light ones can be directly removed and their owned edges reinserted with insertEdges. The heavy ones are removed and their owned edges are returned to be matched in the next round of random settle.
@@ 5. Analysis
We now analyze the cost bounds. We consider both work and 0pt, although the 0pt analysis is straightforward and left until the end. As in prior dynamic maximal matching algorithms (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024), it is useful to consider the lifetime of a match from when it is created until it is deleted. This is referred to as an epoch. It is also useful to distinguish epochs that are deleted by the user in the call to deleteEdges, vs. ones that are deleted internally by the algorithm by a random settle. The first kind are referred to as natural deletions of an epoch and the second as induced deletions of an epoch. In the following we will equate the epoch $e$ with the edge $e$ corresponding to its match.
@@ 5.1. Direct Work: Charging to all Epochs
The expected work of insertEdges(E)(E) is $O(r|E|)$.
@@
The cost of each round of random settle is either covered by matches in the current round or matches in the next round. Specifically, all but raising of matches edges that end up being bloated, and the cost of deleteMatchedEdges at the end, will be covered by the current round.
@@
Putting the work costs together we get the following work bound.
@@
Follows from Lemma 5.2, 5.3, and 5.4, which combined cover all work. ∎
@@ 5.2. Total Work: Charging to Natural Epochs
The input edge set $E$ to randomSettle is the union of the owned cross edges of a set of heavy deleted matches $M_{\mbox{\footnotesize old}}$. Each newly matched edge $m\in M_{\mbox{\footnotesize new}}$ can therefore, for accounting purposes, be associated with the match in $M_{\mbox{\footnotesize old}}$ that previously owned it, and we will use $p(m)$ to indicate this deleted match. Each matched edge $m\in M_{\mbox{\footnotesize new}}$ can contribute up to r1r-1 stolen deletes, one for each of its vertices except at least one incident on $p(m)$. Now each $e\in M_{\mbox{\footnotesize old}}$ can have at most $r$ newly matched edges incident on it (otherwise it would not be a matching). Therefore we can associate r(r1)r(r-1) stolen deletes with each $e\in M_{\mbox{\footnotesize old}}$. Furthermore we also associate the one bloated delete that might have caused each $e\in M_{\mbox{\footnotesize old}}$ to rise, and hence there are fewer than $r^{2}$ deletes associated with deleted edge $e$. This covers all the deletes.
@@
$S_{a}=\sum_{e\in M_{\mbox{\footnotesize new}}}|S(e)|=\sum_{e\in M_{\mbox{\footnotesize old}}}|C(e)|\geq 4r^{2}\sum_{e\in M_{\mbox{\footnotesize old}}}2^{l(e)}>2S_{d}$
We note that having the gap between levels be a constant factor ($\alpha=2$ in our case) instead of proportional to $r$ (as in Assadi and Solomon (Assadi and Solomon, 2021)) is important to the proof. In particular, in the last step of the proof of Lemma 5.6 we lose a factor of $\alpha$. Since we have $\alpha=2$, this brings the coefficient from $4$ to $2$ for us, which is fine. But if $\alpha$ was $\Theta(r)$, then the coefficient would be reduced from $4$ to $\Theta(4/r)$, which is too small to be useful. We could have compensated for $\alpha=\Theta(r)$ by increasing the cutoff for being heavy (in isHeavy) by a factor of $r$, to $\Theta(r^{3}2^{l(e)})$, but then the amortized cost of the overall algorithm would also have increased by a factor of $r$, to be $O(r^{4})$ per edge update. Thus, we save a factor of $r$ by using thinner levels. In AS, reducing the level gap from $\Theta(r)$ to $2$ is not necessary because they assign every new match to some specific deletes. This extra cost consideration in our algorithm comes from “shuffling” the sampled edges during the static matching.
@@
Let $S_{a}$ be the total sample size of all new epochs, $S_{i}$ the total sample size of all induced deletions and $S_{n}$ of natural deletes. From Lemma 5.6 we have that in each round of random settle the number of added samples is at least twice the number of samples from induced deletions. At the end of a batch update, the algorithm runs a insertEdges on the remaining unmatched edges. These are therefore not covered by the sample set of a random settle. Due to the termination condition, however, we know that the sample size of previously matched edges is twice as large as the number of the remaining edges. Let $|\mbox{bloated}|$ be the sample size of bloated remaining edges. Because these edges are bloated, they accrued a factor of at least $2r^{2}$ cross edges after rising. Thus, the sample size of deleted matches from the last round of random settle is at most $\frac{|\mbox{bloated}|}{2r^{2}}\leq\frac{|\mbox{bloated}|}{2}$. Therefore, the total sample space of induced deletions is at most the $\frac{1}{4}S_{a}$ from the final round plus the $\frac{1}{2}S_{a}$ from all other rounds, giving $S_{a}>\frac{4}{3}S_{i}$ within each batch deletion and hence also across all batch deletions. Since we start and end empty, we have that $S_{a}=S_{i}+S_{n}$. Putting these together we get $S_{n}>\frac{1}{3}S_{i}$.
@@ 5.3. Overall Cost Bounds
We now consider the overall work and 0pt bounds. We want to bound the total amount of natural sample space destroyed, so that we can bound the overall work with Lemma 5.7. We use $\Phi$, as introduced in Section 3.1, as a measure of sample space destroyed. We demonstrate that $\Phi$ is well-defined in the wider batch-dynamic setting. At a high level, $\Phi$ still functions properly in the wider algorithm because sampled edges are isolated while their matched edge lives. Then, we use $\Phi$ to bound the natural sample space deleted across the entire run of the algorithm.
Firstly we note that if we have a bound for the work from a start where the graph is empty to an end when it is empty again, this implies a bound that is at least as good for any prefix at least half as dense (in total batch size) as the whole sequence. All runs are a prefix of at least half of an empty-to-empty sequence since we can just delete the edges and the number of final deletes is at most the number of prior operations. Hence we need only consider empty-to-empty runs to asymptotically bound the cost for the prefix.
Let $Q$ be the set of natural epochs. Let $d_{1},d_{2},\ldots,d_{T}$ be the set of deletions by the user across all batches. Suppose that we dynamically assign prices to edges and payments $\Phi$ to edge deletions, as described in Section 3.1, as follows. Whenever we call parallelGreedyMatch, we assign prices to the edges involved. When an edge $e$ is deleted, we pay $\Phi(e)$. Then, if $e$ was sampled and unmatched, we decrement 1 from the price of $e$’s associated matched edge. For periods of time when an edge $e^{\prime}$ has no price, we set $\Phi(e^{\prime})=0$.
If a matched edge gets deleted in the same time step as an unmatched edge in its sample space, we will consider both deletes to be early. Accordingly, if a matched edge and $z$ of its sampled edges are deleted in batch, we will subtract $z$ from the matched edge’s price before paying its price.
@@
Assigning prices, as in Section 3.1, is well-behaved:
  1. (1)
    For all times $t$, $\Phi$ is a function.
  2. (2)
    An edge is sampled iff $\Phi(e)>0$. An edge deletion is early iff the edge is sampled.
  3. (3)
    (Extension of Lemma 3.3): Suppose that the user deletes edge $e$ at time $t$. Then $\mathbb{E}[\Phi(e)]\leq 2$.
@@
We proceed as follows.
  1. (1)
    Note that for an edge $e$, if $e$ has a price at time $t$, then by construction $e$’s matched edge still exists in the graph, so $e$ is sampled at time $t$. Note that a sampled edge is owned by exactly one epoch within the instance of parallelGreedyMatch that sampled it. Because sampled edges do not participate in the algorithm until their matched edge is deleted, this edge will not join the sample space of any other epochs. Therefore, $e$ has at most one price at time $t$. Therefore, $\Phi(e)$ has at most one value. Because we set $\Phi(e)$ to 0 if an edge $e$ has no price (is a cross edge), we have that $\Phi(e)$ will map to exactly one value for all edges $e$. Therefore $\Phi$ is well-defined in the dynamic setting.
  2. (2)
    Consider the early delete of an edge $e$. Recall that $e$’s matched edge is $p(e)$, and that $e$ is in the sample space $S_{p(e)}$. By definition of early, $p(e)$ is still matched. Therefore, $p(e)$’s sample space is still sequestered from the rest of the algorithm, and so $e$ is still in $p(e)$’s sample space. Therefore, $e$ is sampled. Similarly, let $e^{\prime}$ be a sampled edge. Because $e^{\prime}$ is sampled, its matched edge $p(e^{\prime})$ is still alive, so the deletion of $e$ is early.
    Note that the scheme in Section 3.1 never assigns payment 0 to a deletion. We conclude that a deleted edge gives positive payment iff the edge is sampled and iff the deletion is early.
  3. (3)
    Note that $\mathbb{E}[\Phi(e)]$ is equal to $Pr[\Phi(e)>0]\mathbb{E}[\Phi(e)|\Phi(e>0)]+Pr[\Phi(e)=0](0)$, which is no more than $\mathbb{E}[\Phi(e)|\textrm{early}]$, which by Lemma 3.3 is no more than $2$.
Lemma 5.9.
The total natural sample space destroyed is no more than the total payment given by all user deletes: $\sum_{e\in Q}|S_{e}|\leq\sum_{t=1}^{T}\Phi(d_{t})$.
Proof.
For this proof, we look at the algorithm after it has finished running. Note that a natural epoch $e$ is destroyed by a user deletion and contributes $|S_{e}|$ to the total natural sample space destroyed, where $|S_{e}|$ is the original size of $e$’s sample space.
Consider the price associated with this same natural epoch $e$, deleted at time $t$. We proceed in a similar fashion to the proof of Lemma 3.4. Suppose that $e$ started with $|S_{e}|$ sample space, has $y$ sample space at the beginning of time $t$, and has $z$ of its sampled edges also deleted at time $t$, in the same batch. We pay yzy-z for deleting $e$, and the deletions of its sampled edges contribute $(|S_{e}|-y)+z$ payment, for $|S_{e}|$ payment total. Because edges have at most one payment value associated to them, this $|S_{e}|$ payment is only associated with the natural epoch $e$. Therefore, the total payment is equal to the sum of the payments associated with each epoch.
Since for each natural epoch, the sample space lost is no more than the payment, the lemma follows. Note that the total payment can be greater than the total natural sample space lost because unmatched sampled deletes from induced epochs pay but do not destroy natural sample space. ∎
@@
In this proof, we view the algorithm before the run begins. By Lemma 5.9, observe that $\mathbb{E}[\sum_{e\in Q}|S_{e}|]\leq\mathbb{E}[\sum_{t=1}^{T}\Phi(d_{t})]$. By Lemma 5.8, observe that $\sum_{t=1}^{T}\mathbb{E}[\Phi(d_{t})]\leq\sum_{t=1}^{T}2=2T$. By Lemma 5.7, the total work of the algorithm is bounded by
$\sum_{e\in Q}Cr^{3}2^{l(e)}\leq\sum_{e\in Q}2Cr^{3}|S_{e}|\leq 2Cr^{3}(2T)\leq O(r^{3}N),$
where $C$ is the constant hidden by Lemma 5.7. ∎